# On the Expressiveness and Length Generalization of Selective State-Space Models on Regular Languages

### v1

@@ Abstract
Selective state-space models (SSMs) are an emerging alternative to the Transformer, offering the unique advantage of parallel training and sequential inference. Although these models have shown promising performance on a variety of tasks, their formal expressiveness and length generalization properties remain underexplored. In this work, we provide insight into the workings of selective SSMs by analyzing their expressiveness and length generalization performance on regular language tasks, i.e., finite-state automaton (FSA) emulation. We address certain limitations of modern SSM-based architectures by introducing the Selective Dense State-Space Model (SD-SSM), the first selective SSM that exhibits perfect length generalization on a set of various regular language tasks using a single layer. It utilizes a dictionary of dense transition matrices, a softmax selection mechanism that creates a convex combination of dictionary matrices at each time step, and a readout consisting of layer normalization followed by a linear map. We then proceed to evaluate variants of diagonal selective SSMs by considering their empirical performance on commutative and non-commutative automata. We explain the experimental results with theoretical considerations. Our code is available at https://github.com/IBM/selective-dense-state-space-model.
@@ 1 Introduction
In Sec. 3, we introduce the first selective SSM capable of perfect ($\geq$ 99.9%) length generalization on FSA emulation using a single layer. We call this model SD-SSM, the Selective Dense State-Space Model. SD-SSM utilizes a dictionary of dense unstructured transition matrices, a softmax selection mechanism that creates a convex combination of a fixed number of transition matrices at each step, and finally applies a readout consisting of layer normalization followed by a linear map. We identify that a common design choice, the presence of a nonlinear readout, prevents SD-SSM from achieving full accuracy on a challenging FSA emulation task. We compare the model with the standard RNN and the LSTM (Hochreiter and Schmidhuber 1997) in terms of the minimal sample length required to generalize in length and find that SD-SSM exhibits much better length generalization. Moreover, running SD-SSM with a parallel algorithm yields a notable speed-up over its sequential implementation.
In Sec. 4, we take a closer look at selective SSMs with diagonal complex transition matrices. We evaluate them on a set of FSA emulation tasks and analyze the effects of different architectural design choices on their performance on a commutative and non-commutative automaton. We find that perfect in-domain accuracy can be achieved on both automata, but length generalization is significantly worse on the non-commutative one. We explain our experimental results with such systems by demonstrating that, under an assumption on the mapping of FSA to selective SSMs, single-layer selective diagonal SSMs obey a tighter upper bound on expressiveness than the one shown in (Merrill, Petty, and Sabharwal 2024).
@@ 3 Experimental Analysis of SD-SSM on Regular Languages
The investigated tasks aim at tracking the state transitions of different FSAs. The experimental code is based on (Delétang et al. 2023) and (Liu et al. 2023). We evaluate our models on seven different FSAs. Parity, Even Pairs, Cycle, and Arithmetic are taken from (Delétang et al. 2023). We further define three automata based on the Cayley diagrams of different algebraic groups. $C_{2}\times C_{4}$, the direct product of cyclic groups $C_{2}$ and $C_{4}$, is a commutative, solvable group with eight states. $D_{4}$, the dihedral group with eight elements, is a non-commutative, solvable group. $A_{5}$ is the group of even permutations of five elements, a non-solvable group with 60 states. The tasks are described in Appendix A [^1] The appendix is available in the preprint (Terzić et al. 2024)..
@@
We additionally ablate the SD-SSM readout. Apart from the desire for simplicity, the SD-SSM’s readout also emerged from the observation that a more complex readout has detrimental effects on the length generalization. On the Arithmetic task, we conducted extensive experiments in which the linear layer in SD-SSM’s readout was replaced by a standard two-layer MLP with the ReLU non-linearity. The hidden layer of the MLP was configured to consist of 64 units, equal to the state size of the model. We varied the learning rate, the $p$ parameter in $l_{p}$ normalization, and we experimented with two regularization techniques, weight decay and dropout on the intermediate activations of the MLP readout. The best accuracy was achieved by using weight decay. The model achieves an accuracy of $71.9\%$, significantly below $99.9\%$ achieved by the linear readout SD-SSM. Further results are shown in Appendix C.
@@
SD-SSM Can Leverage Parallel Scan Algorithm
@@ 4 A Limitation of Diagonal Selective SSMs
with $W_{Re,Im}^{i}\in\mathbb{R}^{n\times d}$, $W_{Re,Im}^{o}\in\mathbb{R}^{n\times n}$, $W_{i}^{y}\in\mathbb{R}^{2n\times 2n}$, $W_{o}^{y}\in\mathbb{R}^{d\times 2n}$, $B\in\mathbb{C}^{n\times d}$, $\odot$ denoting the element-wise product, $\sigma(\cdot)$ an element-wise activation function, in our case ReLU, ||$\cdot$|| ⋅ | the element-wise absolute value function, and $a\oplus b$ denoting the concatenation of the two vectors $a$ and $b$.
The entries along the diagonal of the presented model’s transition matrices are complex numbers whose real and imaginary parts are generated as functions of the input using a two-layer MLP. They are constrained to be on the complex unit circle by dividing its entries by their magnitudes. This system can potentially be unstable if the imaginary component is zero. However, we did not observe this in our experiments.
@@ 6 Conclusion
In this work, motivated by the inability of a wide range of sequence model to emulate arbitrary automata, we have presented SD-SSM. It utilizes a dictionary of dense transition matrices, combined at each time step using a softmax selection mechanism and linear operator normalization, and a readout which consists of layer normalization followed by a linear map. SD-SSM is the first selective state-space model to achieve perfect length generalization on a diverse set of FSA emulation tasks using a single layer.
We then evaluated more efficient selective SSMs with diagonal complex valued transition matrices on a set of FSA emulation tasks. We observed that they exhibit significantly worse length generalization than their dense counterparts. We probed deeper into this result by investigating their performance on two similar automata which differ in one crucial property: commutativity with respect to the inputs. Our experimental analysis confirms that diagonal selective SSMs exhibit a significantly higher degree of length generalization on the commutative automaton compared to the non-commutative automata. We explain the results by drawing a connection between a general mapping of FSA dynamics onto selective SSM weights and linear system theory. Assuming that the selective SSMs do not implement an unintuitive mapping of FSA dynamics, we observe that they indeed cannot model non-commutative automata.

### v2

@@ Abstract
Selective state-space models (SSMs) are an emerging alternative to the Transformer, offering the unique advantage of parallel training and sequential inference. Although these models have shown promising performance on a variety of tasks, their formal expressiveness and length generalization properties remain underexplored. In this work, we provide insight into the workings of selective SSMs by analyzing their expressiveness and length generalization performance on regular language tasks, i.e., finite-state automaton (FSA) emulation. We address certain limitations of modern SSM-based architectures by introducing the Selective Dense State-Space Model (SD-SSM), the first selective SSM that exhibits perfect length generalization on a set of various regular language tasks using a single layer. It utilizes a dictionary of dense transition matrices, a softmax selection mechanism that creates a convex combination of dictionary matrices at each time step, and a readout consisting of layer normalization followed by a linear map. We then proceed to evaluate variants of diagonal selective SSMs by considering their empirical performance on commutative and non-commutative automata. We explain the experimental results with theoretical considerations.
Codehttps://github.com/IBM/selective-dense-state-space-model
@@ 1 Introduction
In Sec. 3, we introduce the first selective SSM capable of perfect ($\geq$ 99.9%) length generalization on FSA emulation using a single layer. We call this model SD-SSM, the Selective Dense State-Space Model. SD-SSM utilizes a dictionary of dense unstructured transition matrices, a softmax selection mechanism that creates a convex combination of a fixed number of transition matrices at each step, and finally applies a readout consisting of layer normalization followed by a linear map. We identify that a common design choice, the presence of a nonlinear readout, prevents SD-SSM from achieving full accuracy on a challenging FSA emulation task. We compare the model with the standard RNN and the LSTM (Hochreiter and Schmidhuber 1997) in terms of the minimal sample length required to generalize in length and find that SD-SSM exhibits better length generalization. Moreover, running SD-SSM with a parallel algorithm yields a notable speed-up over its sequential implementation.
In Sec. 4, we take a closer look at selective SSMs with diagonal complex transition matrices. We evaluate them on a set of FSA emulation tasks and analyze the effects of different architectural design choices on their performance on a commutative and non-commutative automaton. We find that perfect in-domain accuracy can be achieved on both automata, but length generalization is significantly worse on the non-commutative one. We explain our experimental results with such systems by demonstrating that, under an assumption on the mapping of FSA to selective SSMs, single-layer selective diagonal SSMs are restricted to emulating commutative automata.
@@ 3 Experimental Analysis of SD-SSM on Regular Languages
The investigated tasks involve tracking the state transitions of different FSAs. The experimental code is based on (Delétang et al. 2023) and (Liu et al. 2023). We evaluate our models on seven different FSAs. Parity, Even Pairs, Cycle, and Arithmetic are taken from (Delétang et al. 2023). We further define three automata based on the Cayley diagrams of different algebraic groups. $C_{2}\times C_{4}$, the direct product of cyclic groups $C_{2}$ and $C_{4}$, is a commutative, solvable group with eight states. $D_{4}$, the dihedral group with eight elements, is a non-commutative, solvable group. $A_{5}$ is the group of even permutations of five elements, a non-solvable group with 60 states. The tasks are described in Appendix A [^1] The appendix is available in the preprint (Terzić et al. 2024)..
@@
We additionally ablate the SD-SSM readout. Apart from the desire for simplicity, the SD-SSM’s readout also emerged from the observation that a more complex readout has detrimental effects on the length generalization. On the Arithmetic task, we conducted extensive experiments in which the linear layer in SD-SSM’s readout was replaced by a standard two-layer MLP with the ReLU non-linearity. The hidden layer of the MLP was configured to consist of 64 units, equal to the state size of the model. We varied the learning rate in {2e-5, 1e-4, 5e-4}, the $p$ parameter in $l_{p}$ normalization in {1.1, 1.2, 1.3}, and we experimented with two regularization techniques, weight decay in {0, 1e-4, 1e-3} and dropout in {0.1, 0.2, 0.5} on the intermediate activations of the MLP readout. The best accuracy was achieved by using weight decay. The model achieves an accuracy of $71.9\%$, significantly below $99.9\%$ achieved by the linear readout SD-SSM. Further results are shown in Appendix C.
@@
SD-SSM Can Leverage Parallel Scans
@@ 4 A Limitation of Diagonal Selective SSMs
with $W_{Re,Im}^{i}\in\mathbb{R}^{n\times d}$, $W_{Re,Im}^{o}\in\mathbb{R}^{n\times n}$, $W_{i}^{y}\in\mathbb{R}^{2n\times 2n}$, $W_{o}^{y}\in\mathbb{R}^{d\times 2n}$, $B\in\mathbb{C}^{n\times d}$, $\beta\in[0,1]$, $\odot$ denoting the element-wise product, $\sigma(\cdot)$ an element-wise activation function, in our case ReLU, ||$\cdot$|| ⋅ | the element-wise absolute value function, and $a\oplus b$ denoting the concatenation of the two vectors $a$ and $b$.
The entries along the diagonal of the presented model’s transition matrices are complex numbers whose real and imaginary parts are generated as functions of the input using a two-layer MLP. They are constrained to be on the complex circle of radius $\beta$, where $\beta$ is a hyperparameter.
@@ 6 Conclusion
In this work, motivated by the inability of a wide range of sequence models to emulate arbitrary automata, we have presented SD-SSM. It utilizes a dictionary of dense transition matrices, combined at each time step using a softmax selection mechanism and operator normalization, and a readout which consists of layer normalization followed by a linear map. SD-SSM is the first selective state-space model to achieve perfect length generalization on a diverse set of FSA emulation tasks using a single layer.
We then evaluated more efficient selective SSMs with diagonal complex valued transition matrices on a set of FSA emulation tasks. We observed that they exhibit significantly worse length generalization than their dense counterparts. We probed deeper into this result by investigating their performance on two similar automata which differ in one crucial property: commutativity with respect to the inputs. Our experimental analysis confirms that diagonal selective SSMs exhibit a significantly higher degree of length generalization on the commutative automaton compared to the non-commutative one. We explain the results by drawing a connection between a general mapping of FSA dynamics onto selective SSM weights and linear system theory. Assuming that the selective SSMs do not implement an unintuitive mapping of FSA dynamics, we observe that they indeed cannot model non-commutative automata.